A guage records the pressure of a tire as 160000Nm/2. Assuming the volume is unchange, what does the guage read at 30degrees clesius ? The reading ona Tyre reads the pressure above atmospheric pressure. The atmosphere pressure can be assumed to be 100000Nm/2



1
Expert's answer
2021-01-30T10:42:33-0500
"P_2=P_1\frac{T_2}{T_1}\\P_2^g+P_a=(P_1^g+P_a)\frac{T_2}{T_1}
\\P_2^g+100000=(160000+100000)\frac{273+30}{273}\\P_2^g=190000\ Pa"